Solve the following linear equations
x + 7 – 8x/3 = 17/6 – 5x/2 

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student

x+7–8x/3=17/6–5x/2
⇒x–8x/3 + 5x/2 = 17/6–7
⇒(6x–16x+15x)/6=(17-42)/6
⇒5x/6=-25/6
⇒5x=-25
⇒x=-5

Thanks